INA FILM ACTOR CHARLES COBURN PLAYS AN ELDERLY UNCLE CHARACTER CRITIZIED FOR MARRYING A WOMEN WHEN HE IS 3 TIMES HER AGE. HE WRITTENLY REPLIES AH BUT IN 20 YEARS I SHALL ONLY BE TWICE HER AGE. HOW OLD IS THE UNCLE AND THE WOMEN? SET UP A LINEAR EQUATION TO SOLVE THE PROBLEM.
Answer by Alan3354(69443)

Poorly typed, all CAPS.
---------------
U = Uncle's age
W = Wife's age
---------------
Now:
U = 3W
---
In 20 years:
U + 20 = 2(W + 20)
Sub for U
3W + 20 = 2W + 40
W = 20 (wife's age now)
U = 60 (uncle's age now)